OpenStudy (anonymous):

Solve the proportion. 16/7=4.8/h a. 2.1 c. 3.5 b. 3.2 d. 4.8

OpenStudy (rsadhvika):

cross multiply

OpenStudy (jh3power):

16/7 h=4.8 h=4.8 * 7/16 h=2.1

OpenStudy (rsadhvika):

dont u think its rude to close the quesiton when ppl are working on it trying to help you out @hadleydarcym

OpenStudy (anonymous):

sorry

OpenStudy (rsadhvika):

its okay, hope you figured the solution already

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

other tag please. lol